### v2

# Detection of Multiband Lensed Gravitational Waves from Dark Matter Halos with Deep Learning

Abstract
Lensed gravitational waves (GWs) propagate through perturbed gravitational potentials and produce interference patterns in amplitude and phase. These patterns contain information about dark matter substructures, yet their signatures are often buried in detector noise. To address this challenge, we develop a deep learning framework for identifying lensed signals under multiband observations. The model processes binary neutron star waveforms, with the early inspiral observed by the DECi-hertz Interferometer Gravitational-wave Observatory (DECIGO) and the late inspiral captured by the Einstein Telescope (ET). These two bands approximately correspond to the wave optics and geometric optics regimes. We consider five classes of data: pure noise, unlensed signals, and three lensed signals generated by the Singular Isothermal Sphere, Cored Isothermal Sphere, and Navarro–Frenk–White profiles. Using a dataset of $10^{6}$ samples with signal to noise ratio values from 5 to 100, the model achieves an accuracy of $97.0\%$, notably exceeding the performance of single detector configurations such as DECIGO or ET. The model maintains an accuracy above $63\%$ when the SNR is below $20$, and the accuracy increases to $99.8\%$ once the SNR exceeds $80$. These results demonstrate that multiband observations improve the detectability of lensing effects and provide an efficient and robust approach for automated identification of lensed GWs.
preprint: APS/123-QED
I Introduction
[…]
Recent theoretical and numerical studies have shown that lensed GWs introduce distinctive amplitude modulations and time delays in the frequency domain, providing a unique probe of dark matter halos and primordial black holes [tambalo2023lensing, leung2025wave]. Building on these findings, upcoming third generation detectors such as the Einstein Telescope (ET) and the space based DECi hertz Interferometer Gravitational wave Observatory (DECIGO) are expected to enable multiband observations that further enhance the detectability and parameter estimation precision of lensed GWs [abe2025multi, mukherjee2020multimessenger, li2023targeted]. However, identifying and characterising lensed GWs remain challenging. Traditional matched filtering and Bayesian inference have played a central role in gravitational wave detection and parameter estimation, and continue to provide reliable results across various conditions. As observational sensitivity increases and the volume of detected events grows, these approaches may face practical limitations related to computational efficiency and waveform modeling [janquart2021fast, li2018gravitational]. In particular, low SNR conditions and interference induced modulations can further complicate signal recovery. These developments highlight the importance of exploring complementary and adaptive analysis frameworks for next generation gravitational wave data.
With the rapid development of deep learning and machine learning techniques, these methods have been progressively introduced into gravitational wave data analysis and lensed GWs identification [kim2021identification, goyal2021rapid, li2025identification, liu2025identifying, chan2025detectability, chan2025identification]. Compared with traditional matched filtering and Bayesian inference approaches, deep learning offers advantages in automatic feature extraction and efficient inference. Convolutional neural networks (CNNs) are capable of learning informative representations from time domain or time frequency domain data, which can help mitigate the dependence on predefined waveform templates. Moreover, their inference phase is computationally efficient, making them particularly suitable for real time identification and filtering of the massive number of GWs events expected in future observations [singh2018classifying, goyal2021rapid]. Recent studies have demonstrated the potential of deep learning in identifying lensed GWs [wilde2022detecting], further confirming its applicability under various conditions. The advent of third generation detectors has also introduced new opportunities to overcome current limitations through multiband observations. Space based detectors such as the DECIGO exhibits higher sensitivity to low frequency wave optics interference effects, while ground based detectors such as the ET offer superior resolution for geometric optics magnification and phase delay phenomena at mid to high frequencies [abe2025multi]. This complementarity across frequency bands implies that, by fusing data from multiple detectors, one can simultaneously exploit multiple scale physical features including diffraction and geometric magnification—to mitigate parameter degeneracies and enhance the accuracy of lensed GWs identification. A deep learning based, multiple detector framework for lensed GWs recognition is therefore expected to play a crucial role in future third generation and space based GWs detectors, providing new methodologies for probing dark matter structures and constraining cosmological parameters.
Building on previous studies, this work proposes a deep learning framework for identifying lensed GWs from binary neutron star (BNS) under multiband observations. The framework uses an improved residual convolutional neural network to perform end to end classification of time domain waveforms in detector noise. Two experimental sets were designed. The first used a unified dataset spanning a signal to noise ratio (SNR) range of 5–100 to train models for single detectors ET, DECIGO, and a multiple detector framework (ET+DECIGO), to investigate the performance improvement achieved through multiband observations. The second divided the data into five SNR intervals to examine performance dependence on SNR. These findings demonstrate that multiband observations significantly enhance the detection and classification of lensed GWs in noisy environments, and highlight their potential for precise automated recognition in future multiband observations.
The paper is structured as follows. Section I outlines the scientific motivation and background. Section II introduces the theoretical framework of lensed GWs and data generation. Section III describes the detector configuration and noise modeling. Section IV presents the dataset construction and simulation setup. Section V details the deep learning architecture and training strategy. Section VI reports the results and analyses under different detector configurations and SNR. Section VII concludes with a summary and future outlook.
II Theoretical Framework
This section outlines the theoretical framework and modelling approach for describing the propagation of GWs through dark matter halo lenses. It introduces the propagation equation under the weak field approximation and summarises the principal characteristics of the adopted mass distribution models.
II.1 Fundamental Theory of Lensed GWs
In the weak field limit, the propagation of GWs through an inhomogeneous spacetime can be described as a perturbation of the background Friedmann–Lemaître–Robertson–Walker (FLRW) metric. Denoting the background metric as $g^{B}_{\mu\nu}$, the perturbed metric is written as
$g_{\mu\nu}=g^{B}_{\mu\nu}+a^{2}h_{\mu\nu},$ (1)
where $a$ is the cosmological scale factor and $h_{\mu\nu}$ is the dimensionless tensor perturbation representing the GWs disturbance. The Newtonian potential of the lens, $\Phi(\mathbf{x})$, satisfies $\Phi\ll 1$ to ensure the validity of the linear approximation. Under this condition, the GWs propagation equation can be approximated as a scalar wave equation with a potential term [abe2025multi]:
$\partial_{\mu}\!\left(\!\sqrt{-g^{B}}\,g^{\mu\nu}_{B}\,\partial_{\nu}h\!\right)=0,$ (2)
where $h(\tau,\mathbf{x})$ is the scalar form of the perturbation. Transforming into the frequency domain yields a Helmholtz-like equation [peters1974index]:
$\left(\nabla^{2}+\omega^{2}\right)\tilde{h}=4\omega^{2}\Phi\,\tilde{h}$, (3)
where $\tilde{h}$ is the frequency domain waveform and $\omega=2\pi f$ is the angular frequency.
To quantify the lensing modulation, the complex amplification factor is defined [nakamura1999wave, takahashi2003wave]:
$F\equiv\frac{\tilde{h}}{\tilde{h}_{\mathrm{UL}}},$ (4)
where $\tilde{h}_{\mathrm{UL}}$ denotes the unlensed waveform. Under the thin lens and Born approximations, $F$ can be expressed as a two dimensional integral over the lens plane [nakamura1999wave, takahashi2003wave]:
$F(w,y)=\frac{w}{2\pi i}\!\int\!d^{2}x\,\exp\!\left[iw\,T(\mathbf{x},\mathbf{y})\right],$ (5)
where $w$ is the dimensionless frequency parameter proportional to both the observing frequency and the lens mass, $y$ is the dimensionless impact parameter, and $T(\mathbf{x},\mathbf{y})$ is the dimensionless time delay function. Here $\bm{\xi}$ and $\bm{\eta}$ denote the two dimensional position vectors on the lens plane and the source plane, respectively. The dimensionless variables are defined as
$x=\frac{\xi}{\xi_{0}},\qquad y=\frac{D_{L}}{\xi_{0}D_{S}}\eta,\qquad w=\frac{\xi_{0}^{2}}{D_{\mathrm{eff}}}\omega,$ (6)
where $\xi_{0}$ denotes the characteristic length scale of the lens, and $\omega=2\pi f$ is the angular frequency of the gravitational wave. The effective distance is given by
$D_{\mathrm{eff}}=\frac{D_{L}D_{LS}}{(1+z_{L})D_{S}},$ (7)
where $D_{L}$, $D_{S}$, and $D_{LS}$ are the angular diameter distances from the observer to the lens, from the observer to the source, and from the lens to the source, respectively. The quantity $z_{L}$ denotes the redshift of the lens. To characterize the lens scale, we define the effective redshifted lens mass as $M_{Lz}=(1+z_{L})M_{L}$, and adopt the normalization $\xi_{0}=\sqrt{4GD_{\mathrm{eff}}M_{Lz}}$, so that,
$w=\frac{8\pi G}{c^{3}}\,M_{Lz}\,f.$ (8)
Hence, $w$ reflects the ratio between the Schwarzschild radius of the lens and the wavelength of the gravitational wave. A large $w$ corresponds to the geometrical optics regime ($w\gg 1$), while a small $w$ corresponds to the wave optics regime ($w\ll 1$[abe2025multi]. In this context, DECIGO mainly operates in the wave optics regime due to its lower frequency, whereas ET approaches the geometrical optics regime at higher frequencies.
The time delay function combines the geometric path difference and the gravitational potential:
$T(\mathbf{x},\mathbf{y})=\tfrac{1}{2}|\mathbf{x}-\mathbf{y}|^{2}-\psi(\mathbf{x})-\phi_{m}(\mathbf{y}),$ (9)
where $\psi(\mathbf{x})$ is the projected lensing potential, and the constant $\phi_{m}(\mathbf{y})$ normalises the minimum of $T$ to zero. The potential $\psi(\mathbf{x})$ is linked to the surface mass density $\Sigma(\mathbf{x})$ through the two dimensional Poisson equation:
$\nabla^{2}_{\mathbf{x}}\psi(\mathbf{x})=2\,\frac{\Sigma(\mathbf{x})}{\Sigma_{\mathrm{cr}}},$ (10)
where $\Sigma_{\mathrm{cr}}=[4\pi G(1+z_{L})D_{\mathrm{eff}}]^{-1}$ is the critical surface mass density, and $D_{\mathrm{eff}}$ depends on the source lens observer geometry. The theoretical framework above provide the foundation for the subsequent numerical simulations and data analysis.
II.2 Lens Models
To investigate the lensed amplification effects of dark matter halos on GWs, three spherically symmetric mass density models are considered: the Singular Isothermal Sphere, the Cored Isothermal Spher, and the Navarro–Frenk–White profile. These models represent a gradual transition from dense to diffuse halo structures, spanning typical mass distributions from galactic to cluster scales. To focus on the dominant physical effects and maintain analytical tractability, all lenses are assumed to be perfectly spherical, neglecting secondary perturbations and substructure effects [poon2025galaxy]. The analysis is performed under the thin-lens and weak field approximations, allowing the amplification factor $F(f,y)$ to be derived in a unified formalism.
Singular Isothermal Sphere (SIS)
The SIS model describes self-gravitating systems with constant velocity dispersion. Its three dimensional density is
$\rho(r)=\frac{v^{2}}{2\pi Gr^{2}},$ (11)
where $v$ is the velocity dispersion. The corresponding surface density is
$\Sigma(x)=\frac{v^{2}}{2G\xi_{0}x}.$ (12)
When $v^{2}=G\Sigma_{\mathrm{cr}}\xi_{0}$, the scale $\xi_{0}$ equals the Einstein radius, giving $\kappa(x)=1/(2x)$ and lensing potential
$\psi(x)=x.$ (13)
The frequency domain amplification factor is expressed analytically as a hypergeometric series [matsunaga2006finite]:
$F(w,y)$ $=e^{iwy^{2}/2}\sum_{n=0}^{\infty}\frac{\Gamma(1+n/2)}{n!}\left(2we^{i3\pi/2}\right)^{n/2}$
$\quad\times{}_{1}F_{1}\!\left(1+\frac{n}{2};\,1;\,-\frac{iwy^{2}}{2}\right),$ (14)
where $w$ is the dimensionless frequency parameter. As $y$ decreases, interference fringes become sharper and magnification stronger.
Cored Isothermal Sphere (CIS)
The CIS model extends the SIS profile by introducing a finite core radius $r_{c}$, representing halos with flattened central densities [kormann1994isothermal, flores1996cluster, treu2010strong]. Its three dimensional and surface densities are
$\rho(r)=\rho_{0}\,\frac{r_{c}^{2}}{r^{2}+r_{c}^{2}},\qquad\Sigma(x)=\frac{\pi\rho_{0}x_{c}\xi_{0}}{\sqrt{(x/x_{c})^{2}+1}},$ (15)
where $x_{c}=r_{c}/\xi_{0}$. For $\rho_{0}=\Sigma_{\mathrm{cr}}/(2\pi\xi_{0}x_{c}^{2})$, $\xi_{0}$ corresponds to the Einstein radius. The lensing potential is
$\psi(x)=\sqrt{x^{2}+x_{c}^{2}}+x_{c}\ln\!\left(\frac{2x_{c}}{\sqrt{x^{2}+x_{c}^{2}}+x_{c}}\right).$ (16)
Compared with the SIS model, the CIS profile yields a smoother central potential. Increasing $x_{c}$ weakens focusing and reduces fringe contrast. The amplification factor must be obtained numerically, depending on $(w,y,x_{c})$.
Navarro–Frenk–White (NFW)
The NFW profile, derived from cosmological $N$ body simulations, describes cold dark matter halos [navarro1996structure, navarro1997universal]:
$\rho(r)=\frac{\rho_{s}}{(r/r_{s})(1+r/r_{s})^{2}}.$ (17)
Defining $x_{s}=r_{s}/\xi_{0}$, the surface density is
$\Sigma(x)=2\rho_{s}r_{s}\,\frac{1-g(x/x_{s})}{(x/x_{s})^{2}-1},$ (18)
where
$g(a)=\begin{cases}\dfrac{\arctan\sqrt{a^{2}-1}}{\sqrt{a^{2}-1}},&a>1,\\[6.0pt] \dfrac{\mathrm{arctanh}\sqrt{1-a^{2}}}{\sqrt{1-a^{2}}},&a<1,\\ 1,&a=1.\end{cases}$ (19)
For $\rho_{s}=\Sigma_{\mathrm{cr}}/(4x_{s}^{3}\xi_{0})$, the lensing potential becomes
$\psi(x)=\tfrac{1}{2}\!\left[\ln^{2}\!\left(\frac{x}{2x_{s}}\right)+\big((x/x_{s})^{2}-1\big)g^{2}(x/x_{s})\right].$ (20)
The frequency domain properties of the NFW model are sensitive to $r_{s}$: smaller $r_{s}$ yields more concentrated halos, stronger magnification, denser fringes, and higher modulation contrast.
In this work, the amplification factors $F(w,y)$ for the three representative dark matter halo lens models (SIS, CIS, and NFW)were computed using the GLoW_public library [GLoW_public].
III Detector Configuration and Noise Modeling
This section outlines the detector configurations and noise models adopted in this study. It first introduces the representative detector systems and their sensitivity characteristics, followed by discussions of the antenna response functions (Sec. III.2) and the modelling of detector noise (Sec. III.3). A unified specification of these components ensures geometric and noise consistency between the DECIGO and the ET, establishing a consistent basis for waveform simulation and signal analysis of lensed GWs.
III.1 Detector Systems and Sensitivity Characteristics
[…]
Figure 1: Amplitude spectral densitie (ASD) of DECIGO and ET compared with that of a representative BNS gravitational wave.
III.2 Detector Response Functions
In multiband observations, the detector response functions describe how the detected strain couples to the source direction, polarization, and detector orientation. For an arbitrary GW source, the measured strain at the detector is expressed as
$h_{\mathrm{det}}(t)=F_{+}(t)\,h_{+}(t)+F_{\times}(t)\,h_{\times}(t),$ (21)
where $F_{+}$ and $F_{\times}$ are the antenna pattern functions for the “++” and $\times$” polarization modes, respectively, and $h_{+}$ and $h_{\times}$ denote the corresponding waveform components. These response functions encode the time dependent coupling between the incident direction $(\theta,\phi)$, the polarization angle $\psi$, and the detector geometry.
DECIGO is a space based interferometer with a triangular three arm configuration whose antenna response varies periodically with time. Let $(\bar{\theta}_{S},\bar{\phi}_{S})$ denote the source position in the heliocentric coordinate system and $(\hat{\theta}_{L},\hat{\phi}_{L})$ the orientation of the orbital angular momentum. The observed strain in the detector frame is [yagi2011detector]
$h(t)=F_{+}(\bar{\theta}_{S},\bar{\phi}_{S},\hat{\theta}_{L},\hat{\phi}_{L})\,h_{+}(t)+F_{\times}(\bar{\theta}_{S},\bar{\phi}_{S},\hat{\theta}_{L},\hat{\phi}_{L})\,h_{\times}(t),$ (22)
where
$F_{+}(\theta_{S},\phi_{S},\psi_{S})$ $=\frac{1}{2}\!\left(1+\cos^{2}\theta_{S}\right)\cos(2\phi_{S})\cos(2\psi_{S})$
$\quad-\,\cos\theta_{S}\sin(2\phi_{S})\sin(2\psi_{S}),$ (23)
$F_{\times}(\theta_{S},\phi_{S},\psi_{S})$ $=\frac{1}{2}\!\left(1+\cos^{2}\theta_{S}\right)\cos(2\phi_{S})\sin(2\psi_{S})$
$\quad+\,\cos\theta_{S}\sin(2\phi_{S})\cos(2\psi_{S}).$ (24)
The transformation between the angular parameters $(\theta_{S},\phi_{S},\psi_{S})$ and the heliocentric coordinates is given by
$\theta_{S}(t)$ $=\cos^{-1}\!\left[\frac{1}{2}\cos\bar{\theta}_{S}-\frac{\sqrt{3}}{2}\sin\bar{\theta}_{S}\cos\!\bigl(\hat{\phi}(t)-\bar{\phi}_{S}\bigr)\right],$ (25)
$\phi_{S}(t)$ $=\frac{\pi}{12}+\tan^{-1}\!\left[\frac{\sqrt{3}\cos\bar{\theta}_{S}+\sin\bar{\theta}_{S}\cos\!\bigl(\hat{\phi}(t)-\bar{\phi}_{S}\bigr)}{2\sin\bar{\theta}_{S}\sin\!\bigl(\hat{\phi}(t)-\bar{\phi}_{S}\bigr)}\right],$ (26)
$\psi_{S}$ $=\tan^{-1}\!\left(\frac{a}{b}\right),$ (27)
where
$a$ $=\frac{1}{2}\cos\hat{\theta}_{L}-\frac{\sqrt{3}}{2}\sin\hat{\theta}_{L}\cos[\hat{\phi}(t)-\hat{\phi}_{L}]$
$\quad-\cos\hat{\theta}_{L}\cos^{2}\bar{\theta}_{S}-\sin\hat{\theta}_{L}\sin\bar{\theta}_{S}\cos(\hat{\phi}_{L}-\bar{\phi}_{S}),$ (28)
$b$ $=\frac{1}{2}\sin\hat{\theta}_{L}\sin\bar{\theta}_{S}\sin(\hat{\phi}_{L}-\bar{\phi}_{S})$
$\quad-\frac{\sqrt{3}}{2}\cos\hat{\phi}(t)[\cos\hat{\theta}_{L}\sin\bar{\theta}_{S}\sin\bar{\phi}_{S}-\cos\bar{\theta}_{S}\sin\hat{\theta}_{L}\sin\hat{\phi}_{L}]$
$\quad-\frac{\sqrt{3}}{2}\sin\hat{\phi}(t)[\cos\bar{\theta}_{S}\sin\hat{\theta}_{L}\cos\hat{\phi}_{L}-\cos\hat{\theta}_{L}\sin\bar{\theta}_{S}\cos\bar{\phi}_{S}],$ (29)
and
$\bar{\theta}(t)=\pi/2,\qquad\hat{\phi}(t)=2\pi t/T+c_{0},$ (30)
where $T=1~\mathrm{yr}$ is the orbital period and $c_{0}=0$ is the initial orbital phase.
The Einstein Telescope consists of three co-located interferometers arranged in an equilateral triangle, each having equal arm lengths and an opening angle of $\gamma=\pi/3$. Let $(\theta,\phi)$ denote the source sky location in the detector frame and $\psi$ the polarization angle. For a single interferometer with arm orientation offset $\Delta$, the antenna pattern functions for the plus and cross polarizations are given by [hild2011sensitivity]
$F_{+}(\theta,\phi,\psi;\gamma,\Delta)$ $=-\frac{\sin\gamma}{2}\!\Bigl[(1+\cos^{2}\theta)\sin(2\phi^{\prime})\cos(2\psi)$
+ $\quad+\,2\cos\theta\cos(2\phi^{\prime})\sin(2\psi)\Bigr],$ (31)
$F_{\times}(\theta,\phi,\psi;\gamma,\Delta)$ $=+\frac{\sin\gamma}{2}\!\Bigl[(1+\cos^{2}\theta)\sin(2\phi^{\prime})\sin(2\psi)$
 $\quad-\,2\cos\theta\cos(2\phi^{\prime})\cos(2\psi)\Bigr],$ (32)
where $\phi^{\prime}=\phi+\Delta$. In the ET-D configuration, the three interferometers are oriented with relative rotation angles $\Delta=0$, $2\pi/3$, and $2\pi/3$, corresponding to three $60^{\circ}$ separated arms forming an equilateral triangle layout underground. Each interferometer therefore provides an independent response following the same functional form but rotated in azimuth by $\Delta$.
For data generation, the directional response is averaged over random sky positions and polarization angles, approximating an isotropic source distribution. To ensure geometric consistency between DECIGO and ET for the same GW source, a reference epoch $t_{\mathrm{ref}}$ in the heliocentric coordinate system is defined, and the polarization angle $\psi$ is fixed accordingly using the DECIGO transformation (equations (28)–(29)) with $t=t_{\mathrm{ref}}$:
$\bar{\phi}_{t}=\frac{2\pi}{T}\,t_{\rm ref}+c_{0},\qquad\psi=\tan^{-1}\!\left(\frac{a}{b}\right),$ (33)
In the ET response, the parameters are then set to
$\theta=\bar{\theta}_{S},\qquad\phi=\bar{\phi}_{S},\qquad\psi=\psi(t_{\rm ref}),$ (34)
where $(\bar{\theta}_{S},\bar{\phi}_{S})$ are the source sky coordinates, and $\psi(t_{\mathrm{ref}})$ is the polarization angle determined by the DECIGO geometry at the reference epoch. Here the reference time is set to $t_{\mathrm{ref}}=0$, corresponding to the initial orbital phase of DECIGO. This configuration guarantees that both detectors share identical source direction and polarization orientation when observing the same GW event, ensuring full geometric consistency.
III.3 Noise Modeling and Detector Noise Generation
Based on the one sided noise PSD, we could obtain the time domain noise signal from the one sided PSD; in this paper, we utilized the Python function pycbc.noise.gaussian.noise_from_psd [alex_nitz_2024_10473621], which takes a PSD as input and returns colored Gaussian noise, to simulate the time domain noise received by both the DECIGO and ET detectors. The corresponding PSDs of ET and DECIGO were adopted according to their official sensitivity curves [yagi2011detector, hild2011sensitivity], ensuring that the generated noise realizations accurately reflect the instrumental characteristics across their respective frequency bands.
IV Dataset Construction and Simulation Setup
Under the thin lens and weak field approximations, the lensed GWs can be modelled as a frequency domain modulation of the unlensed (UL) signal. The UL polarization components $\tilde{h}_{+}^{\rm UL}(f)$ and $\tilde{h}_{\times}^{\rm UL}(f)$ are first generated, and the corresponding lensed signals are obtained by multiplying them by the amplification factor $F(w,y)$:
$\tilde{h}_{+}(f)=F(w,y)\,\tilde{h}_{+}^{\rm UL}(f),\qquad\tilde{h}_{\times}(f)=F(w,y)\,\tilde{h}_{\times}^{\rm UL}(f),$ (35)
where $\tilde{h}_{+}(f)$ and $\tilde{h}_{\times}(f)$ denote the “++” and $\times$” polarization components after lensing modulation, and $\tilde{h}_{+}^{\rm UL}(f)$ and $\tilde{h}_{\times}^{\rm UL}(f)$ are their unlensed counterparts. The amplification factor $F(w,y)$ depends on the adopted lens model and includes three classes in this study:
$F(w,y)\in\{F_{\rm SIS}(w,y),\,F_{\rm CIS}(w,y),\,F_{\rm NFW}(w,y)\},$ (36)
corresponding to the SIS, CIS, and NFW dark matter halo models.
Applying an inverse Fourier transform to the modulated frequency domain signals yields the time domain polarization waveforms $h_{+}(t)$ and $h_{\times}(t)$. Considering the detector antenna responses, the measured strain is given by
$h(t)=F_{+}(t;\Omega,\psi)\,h_{+}(t)+F_{\times}(t;\Omega,\psi)\,h_{\times}(t),$ (37)
where $F_{+}$ and $F_{\times}$ are the response functions to the two polarizations, $\Omega$ denotes the sky location of the source, and $\psi$ is the polarization angle. Finally, detector noise $n(t)$ consistent with the detector PSD $S_{n}(f)$ is added to obtain the simulated observation:
$s(t)=h(t)+n(t),$ (38)
where $s(t)$ is the final noisy observation and $n(t)$ is a Gaussian noise sequence with spectral structure defined by $S_{n}(f)$. This procedure comprises four stages, including lensing modulation, Fourier transformation, detector response mapping, and noise injection.
(a) Frequency domain waveforms of DECIGO.
(b) Time domain waveforms of DECIGO.
Figure 2: Frequency and time domain waveforms of BNS ($m_{1}=m_{2}=1.4\,M_{\odot}$, $d_{L}=100~\mathrm{Mpc}$) under SIS, CIS, and NFW lensing models, as observed by DECIGO. The lens parameters are $M_{L}=2\times 10^{3}\,M_{\odot}$, $\psi_{0}=1.0$, $r_{c}=0.3$, $r_{s}=0.3$, and $y=0.3$.
(a) Frequency domain waveforms of ET.
(b) Time domain waveforms of ET.
Figure 3: Frequency and time domain waveforms of ($m_{1}=m_{2}=1.4\,M_{\odot}$, $d_{L}=100~\mathrm{Mpc}$) under SIS, CIS, and NFW lensing models, as observed by ET. The lens parameters are $M_{L}=2\times 10^{3}\,M_{\odot}$, $\psi_{0}=1.0$, $r_{c}=0.3$, $r_{s}=0.3$, and $y=0.3$.
Table 1: Sampling ranges of BNS waveform parameters. The source parameters were fixed at inclination $\iota=1.0$, initial phase $\phi_{0}=0$, and dimensionless spins $\chi_{1}=\chi_{2}=0.1$. The angular parameters $(\bar{\theta}_{S},\bar{\phi}_{S},t_{\mathrm{ref}})$ correspond to the DECIGO response geometry, where the polarization angle $\psi$ is computed at the fixed reference epoch $t_{\mathrm{ref}}=0$ according to Eq. (33). For the ET response, the parameters are determined as $(\theta,\phi,\psi)_{\mathrm{ET}}=(\bar{\theta}_{S},\bar{\phi}_{S},\psi)$ to ensure geometric consistency with DECIGO.
Parameter Symbol Range
Primary mass $m_{1}$ $\mathcal{U}(1,3)~M_{\odot}$
Secondary mass $m_{2}$ $\mathcal{U}(1,3)~M_{\odot}$
Luminosity distance $D_{L}$ $\mathcal{U}(0,10^{3})~\mathrm{Mpc}$
Ecliptic latitude $\bar{\theta}_{S}$ $\cos^{-1}\!\bigl(\mathcal{U}(-1,1)\bigr)$
Ecliptic longitude $\bar{\phi}_{S}$ $\mathcal{U}(0,2\pi)$
Reference epoch $t_{\mathrm{ref}}$ 0 (fixed)
For both the space based DECIGO and the ground based ET detectors, five classes of samples were generated: pure noise (PN), unlensed (UL), and three lensed cases (SIS, CIS, NFW). The physical parameters of the BNS, including component masses and luminosity distances, were randomly and uniformly sampled according to Table 1. To evaluate the impact of SNR on classification performance, the data were stratified by the SNR of the UL signal in noise, divided into five intervals: 552020, 20204040, 40406060, 60608080, and 8080100100. Each SNR interval contained an independent dataset comprising five signal classes (PN, UL, SIS, CIS, NFW) with $4\times 10^{4}$ samples per class, yielding $2\times 10^{5}$ samples per set. All signals shared the same parameter sampling scheme but differed in noise realizations, lens classes, and modulation parameters. For each detector, datasets were generated following identical SNR intervals and sample counts, ensuring direct comparability between DECIGO and ET under matched physical conditions. The definitions of all sample classes are listed in Table 2.
After generating the SNR stratified datasets, the data were further organized for different experimental configurations. For each detector (ET or DECIGO), all SNR intervals were merged into a unified full range dataset spanning 55100100, comprising roughly $10^{6}$ samples, to train single detector models and evaluate overall classification performance. To enable a fair comparison with models trained on individual detectors, a model was trained on the combined dataset from the multiple detector configuration (ET+DECIGO), within the same SNR range. This approach allows us to evaluate the potential performance gains introduced by multiband observations. Additionally, to investigate the effect of SNR, we trained and evaluated separate models across five distinct SNR intervals.
[…]
For the adopted parameters, the characteristic time delay between multiple images can be estimated as[abe2025multi]
$\Delta t_{\mathrm{d}}\simeq 0.03\left(\frac{M_{Lz}}{3\times 10^{3}M_{\odot}}\right)\left(\frac{y}{0.3}\right)~\mathrm{s}.$ (39)
We found that in the ET frequency band, the inspiral waveform of the BNS system yields a time delay close to zero under these conditions. Therefore, in the subsequent analysis, the interference between multiple images was neglected. Representative frequency and time domain waveforms generated under these settings are illustrated in Fig. 2 and Fig. 3, corresponding to a binary neutron star system with the same lens parameters.
Detector noise was generated based on the PSD specific to each detector. After applying lensing modulation, detector response, and noise addition, the time domain signals for each detector were individually adjusted in length. Longer waveforms were truncated, while shorter ones were padded with zeros. This ensured consistent alignment in time within each detector’s dataset. The resulting dataset includes standardized waveforms across five classes, multiple SNR intervals, and both single and multiple detector configurations.
Table 2: frequency domain construction and time domain forms of the five sample classes (each $\sim 10^{5}$ samples). Here, $\tilde{h}^{\rm UL}(f)$ denotes the UL frequency domain waveform, $F(w)$ is the frequency domain amplification factor, $h(t)$ is the detector response convolved strain, and $n(t)$ is detector noise.
Class frequency domain construction time domain form
PN - $n(t)$
UL $\tilde{h}^{\rm UL}(f)$ $h(t)+n(t)$
SIS lens $F_{\rm SIS}(w,y)\,\tilde{h}^{\rm UL}(f)$ $h^{\rm SIS}(t)+n(t)$
CIS lens $F_{\rm CIS}(w,y)\,\tilde{h}^{\rm UL}(f)$ $h^{\rm CIS}(t)+n(t)$
NFW lens $F_{\rm NFW}(w,y)\,\tilde{h}^{\rm UL}(f)$ $h^{\rm NFW}(t)+n(t)$
Figure 4 shows the SNR distributions of all sample classes in the constructed dataset. DECIGO and ET exhibit complementary sensitivities: DECIGO is more responsive to low frequency, interference dominated features, whereas ET is more sensitive in the mid to high frequency regime dominated by geometric amplification.
(a) SNR distribution for DECIGO.
(b) SNR distribution for ET.
Figure 4: SNR distributions of simulated samples for DECIGO and ET detectors.
V Deep Learning Architecture and Training Strategy
To enable efficient and robust classification of dark matter halo lensing effects, we developed a residual stacked one dimensional convolutional neural network (ResNet-like 1D CNN) that operates directly on time domain waveform inputs. The model classifies five classes: PN, UL signals, and three lensed cases corresponding to the SIS, CIS, and NFW models. By processing standardized time domain strain sequences without frequency domain transformation or handcrafted feature extraction, the network retains temporal continuity and local interference structures, which are expected to contribute to better recognition of weak modulations.
The architecture comprises multiple residual convolutional modules, each containing two convolution–normalization–activation (ReLU) layers with identity shortcuts to facilitate feature propagation and alleviate vanishing gradient or degradation effects. The input first passes through an initial 1D convolutional layer for primary feature extraction, followed by five stacked residual blocks forming deep hierarchical channels. Each block applies a $3\times 1$ convolution kernel, Batch Normalization, and a ReLU activation. A global average pooling layer compresses the temporal dimension, and a fully connected layer with a Softmax activation outputs the five class probability distribution. The complete single branch network functions as an end to end multiple class classifier (Fig. 5).
Figure 5: Architecture of the residual stacked convolutional network.
To evaluate the benefit of multiband observations, two network configurations were implemented. The first is a single branch network processing input waveforms from either DECIGO or ET individually. The second is a dual branch fusion network designed for multiple detector inputs, where each branch independently processes data from one detector using identical ResNet substructures. Deep feature maps from both branches are concatenated in feature space and fused via a fully connected layer to form a unified representation. This late fusion approach preserves detector specific feature learning while leveraging the complementarity between the low frequency (DECIGO) and mid to high frequency (ET) bands, thereby enhancing discrimination of both interference–diffraction patterns and geometric magnification effects (Fig. 6).
Figure 6: A dual branch network that combines features from two detectors (ET and DECIGO) using late fusion.
All input waveforms were standardized to zero mean and unit variance to improve numerical stability and convergence. The model was optimized using sparse categorical cross entropy loss and the Adam optimizer with an initial learning rate of $10^{-3}$. The ReduceLROnPlateau strategy was employed to decrease the learning rate adaptively when the validation loss stagnated, and EarlyStopping was used to terminate training when no improvement was observed for several epochs. Training was performed for up to 200 epochs with a batch size of 128 and a 10% validation split. All experiments used fixed length waveform sequences and five classes output labels. Model training and inference were accelerated using an NVIDIA A800 80G GPU and two Intel(R) Xeon(R) Silver 8480 CPU.
VI Results and Analysis
This section evaluates the model’s classification performance under different detector setups and SNR conditions. The analysis is divided into two parts. First, using datasets that span the full SNR range of unlensed signals (0100100), we evaluate and compare the classification results of models trained on individual detectors (ET and DECIGO) with those trained on combined data from both detectors. This comparison helps to quantify the effect of using multiple detectors on overall performance. Second, we examine how classification accuracy varies with SNR within the multiple detector configuration, to better understand how signal strength influences detection sensitivity and the reliability of lens classification.
VI.1 Single and Multiple Detector Performance
Using a comprehensive dataset spanning the 0100100 SNR range, we evaluated the model’s classification performance across five classes: PN, UL signals, and three representative lensing models (SIS, CIS, and NFW). Three input configurations were tested: single detector ET, single detector DECIGO, and the multiple detectors (ET+DECIGO). The dataset was divided into training and testing subsets in a 9:19{:}1 ratio. Each class contained $4\times 10^{4}$ samples, corresponding to $3.6\times 10^{4}$ training and $4\times 10^{3}$ testing samples per class, giving a total of $1.8\times 10^{5}$ training and $2\times 10^{4}$ testing samples.
Model performance was evaluated using Accuracy, Precision, Recall, $F_{1}$ score, and Macro Average metrics. Accuracy quantifies the overall classification correctness and is defined as
$\mathrm{Accuracy}=\frac{TP+TN}{TP+TN+FP+FN},$ (40)
where $TP$, TNTN, FPFP, and $FN$ denote true positives, true negatives, false positives, and false negatives, respectively. Although the above definition corresponds to binary classification, the accuracy reported in this study represents the overall mean value computed across all five lensing classes, providing a single metric that summarizes the model’s global classification performance. Precision and Recall measure the reliability and completeness of the predictions:
$\mathrm{Precision}=\frac{TP}{TP+FP},\qquad\mathrm{Recall}=\frac{TP}{TP+FN}.$ (41)
The $F_{1}$ score provides a harmonic mean of Precision and Recall:
$F_{1}=\frac{2\,\mathrm{Precision}\times\mathrm{Recall}}{\mathrm{Precision}+\mathrm{Recall}}.$ (42)
For multiple class classification, Macro Averaging is applied to balance contributions among all $C$ classes:
$\mathrm{Micro~average}(M)=\frac{1}{C}\sum_{i=1}^{C}M_{i},$ (43)
where $M_{i}$ is the metric for class $i$.
To further analyse inter class discrimination, confusion matrices were plotted to visualise misclassification relationships among classes, and Receiver Operating Characteristic (ROC) curves were constructed to evaluate the trade off between the True Positive Rate (TPR) and False Positive Rate (FPR). The Area Under the Curve (AUC) quantifies the model’s discriminative ability:
$\mathrm{AUC}=\int_{0}^{1}\mathrm{TPR}(\mathrm{FPR})\,d(\mathrm{FPR}),$ (44)
where larger AUC values (approaching 1) indicate stronger separation capability.
Comprehensive comparisons under different input configurations are summarised in Tables 35. The accuracy, average $F_{1}$, and average AUC trends are shown in Fig. 7, and the corresponding ROC curves and confusion matrices are presented in Fig. 8 and Fig. 9, respectively.
Figure 7: The classification Accuracy, average $F_{1}$, and average AUC for ET, DECIGO, and joint (ET+DECIGO) models.
Table 3: Classification performance metrics for the ET model.
Class Precision Recall $F_{1}$ score
PN 0.7528 0.7275 0.7400
UL 0.4347 0.7205 0.5423
SIS 0.6437 0.5999 0.6210
CIS 0.7147 0.6178 0.6627
NFW 0.7752 0.4494 0.5689
Macro average 0.6642 0.6230 0.6270
Accuracy 0.6230
Table 4: Classification performance metrics for the DECIGO model.
Class Precision Recall $F_{1}$ score
PN 0.9993 1.0000 0.9996
UL 0.9294 0.9945 0.9609
SIS 0.5183 0.5626 0.5396
CIS 0.9301 0.4149 0.5738
NFW 0.4766 0.6636 0.5548
Macro average 0.7707 0.7271 0.7257
Accuracy 0.7280
Table 5: Classification performance metrics for the joint (ET+DECIGO) model.
Class Precision Recall $F_{1}$ score
PN 1.0000 1.0000 1.0000
UL 0.9960 0.9998 0.9979
SIS 0.9441 0.9068 0.9251
CIS 0.9146 0.9987 0.9548
NFW 1.0000 0.9444 0.9714
Macro average 0.9709 0.9699 0.9698
Accuracy 0.9700
The ET model demonstrates limited classification capability in the mid to high frequency regime, achieving an accuracy of $0.623$ and a macro averaged $F_{1}$ score of $0.627$. This indicates that high frequency information alone is insufficient to capture the wave optics features of lensed GWs signals. For the “PN” and “UL” classes, the Precision and Recall values are $0.75/0.73$ and $0.43/0.72$, respectively, revealing substantial false negatives and misclassifications under low SNR conditions. Among the lensed classes, the model’s recognition remains limited, with $F_{1,\text{SIS}}=0.621$, $F_{1,\text{CIS}}=0.663$, and $F_{1,\text{NFW}}=0.569$. In particular, the NFW class exhibits a Recall of only $0.449$, reflecting poor sensitivity to weak interference structures. The generally low $F_{1}$ scores suggest that the model struggles to maintain a balance between precision and recall, implying incomplete extraction of lensing related features.
The confusion matrix (Fig. 9(a)) further reveals notable confusion between the SIS and CIS classes, with roughly $22\%$ of samples misclassified, and several NFW signals incorrectly labelled as UL or PN. The ROC curve (Fig. 8) shows that the ET model’s TPR remains low in the small FPR region, indicating weak detection performance under strict classification thresholds. This degradation arises primarily because the network fails to capture low frequency phase interference patterns, leading the decision boundaries to depend largely on amplitude variations. As a result, feature separability is reduced, producing unstable classification outcomes. Nevertheless, the ET model retains moderate sensitivity to strong geometric amplification signatures from high mass lenses, providing complementary information for subsequent multiband fusion analysis.
(a) micro average
(b) PN
(c) UL
(d) SIS
(e) CIS
(f) NFW
Figure 8: Comparison of ROC curves across five classification tasks: (a) micro average, (b) PN, (c) UL, (d) SIS, (e) CIS, and (f) NFW.
The DECIGO model, leveraging low frequency diffraction and phase interference information, achieves a notable improvement in overall performance. Its accuracy reaches $0.728$, with a macro averaged $F_{1}$ score of $0.726$, demonstrating enhanced sensitivity to interference structures in the low frequency band. For the “PN” and “UL” classes, the $F_{1}$ scores reach $0.9996$ and $0.961$, respectively, indicating near perfect recognition. However, for the three lensing cases (SIS, CIS, and NFW), the $F_{1}$ scores are $0.540$, $0.574$, and $0.555$, respectively, revealing residual confusion. The confusion matrix (Fig. 9(b)) shows that misclassification between SIS and CIS occurs at a rate of approximately 25–30%, while around 72% of NFW signals are incorrectly labelled as other lens classes.
ROC analysis (Fig. 8) indicates that DECIGO maintains a high TPR even at low FPR, with an average AUC of about $0.90$, substantially outperforming ET. This result confirms that low frequency data enhances both phase sensitivity and noise robustness. Overall, DECIGO achieves superior classification stability in weak signal regimes compared with ET, demonstrating the advantage of low frequency interference patterns under complex noise conditions.
(a) ET
(b) DECIGO
(c) ET+DECIGO
Figure 9: Comparison of confusion matrices for three configurations: (a) ET, (b) DECIGO, and (c) joint (ET+DECIGO).
The joint model (ET+DECIGO) demonstrates a substantial performance gain through multiband fusion. It achieves an accuracy of $0.970$, with macro averaged Precision, Recall, and $F_{1}$ values of $0.971$, $0.970$, and $0.970$, respectively (Table 5), clearly surpassing both single detector models. For the “PN” and “UL” classes, $F_{1}$ scores reach $1.000$ and $0.998$, indicating near perfect recognition. The three lensed classes also show significant improvement, with $F_{1,\text{SIS}}=0.925$, $F_{1,\text{CIS}}=0.955$, and $F_{1,\text{NFW}}=0.971$.
The confusion matrix (Fig. 9(c)) exhibits strong diagonal dominance, with mutual misclassification between SIS and CIS below $2\%$, and NFW misclassification below $1\%$. The ROC curves (Fig. 8) show that the joint model completely encompasses both single detector models, with a micro average AUC approaching unity and individual class AUCs improving by about 0.100.100.250.25. This demonstrates a clear enhancement in both sensitivity and stability.
The accuracy trend (Fig. 7) further confirms the monotonic improvement from ET ($62.3\%$) to DECIGO ($72.8\%$) to the joint model ($97.0\%$). The superiority of the joint configuration arises from multiband feature complementarity: the low frequency channel contributes phase interference information, while the high frequency channel captures amplitude and propagation delay signatures. Their combination enhances the model’s discriminative capability and noise robustness.
Overall, based on the metrics in Tables 35 and the visual comparisons in Figs. 89, the ET+DECIGO joint model outperforms both single detector counterparts in accuracy, confusion structure, and robustness, confirming the physical and statistical advantages of multiband data fusion for detecting lensed GWs in noisy environments.
VI.2 Performance of the Multiple Detector at Different SNR
To evaluate the robustness and sensitivity of the multiple detector (ET+DECIGO) model under varying SNR, the model was trained and evaluated separately within five SNR intervals: 552020, 20204040, 40406060, 60608080, and 8080100100. Each interval included five classes (PN, UL, SIS, CIS, NFW) with a total of $2\times 10^{5}$ samples ($4\times 10^{4}$ per class). The data were divided into training and testing subsets with a 9:19{:}1 split. Model performance was quantified using Accuracy, Precision, Recall, $F_{1}$ score, and Macro Average metrics. Additionally, confusion matrices (Fig. 12) and ROC curves (Fig. 11) were plotted to visualise class separation across different SNR regimes. Quantitative metrics for each interval are summarised in Tables 610, and the evolution of accuracy, average $F_{1}$, and average AUC with increasing SNR is shown in Fig. 10.
Figure 10: The classification Accuracy, average $F_{1}$, and average AUC for the joint (ET+DECIGO) model across five SNR intervals.
Table 6: Classification metrics of the joint (ET+DECIGO) model in the SNR range 5–20.
Class Precision Recall $F_{1}$ score
PN 1.0000 1.0000 1.0000
UL 0.7791 0.8316 0.8045
SIS 0.5114 0.3910 0.4431
CIS 0.5662 0.1460 0.2321
NFW 0.4214 0.7901 0.5496
Macro average 0.6556 0.6317 0.6059
Accuracy 0.6339
Table 7: Classification metrics of the joint (ET+DECIGO) model in the SNR range 20–40.
Class Precision Recall $F_{1}$ score
PN 1.0000 0.9975 0.9987
UL 0.9081 0.9792 0.9423
SIS 0.5473 0.6379 0.5891
CIS 0.8431 0.0109 0.0215
NFW 0.5239 0.8991 0.6620
Macro average 0.7645 0.7049 0.6427
Accuracy 0.7074
Table 8: Classification metrics of the joint (ET+DECIGO) model in the SNR range 40–60.
Class Precision Recall $F_{1}$ score
PN 1.0000 1.0000 1.0000
UL 0.9997 1.0000 0.9999
SIS 0.8472 0.9992 0.9170
CIS 1.0000 0.8196 0.9009
NFW 0.9993 0.9998 0.9995
Macro average 0.9692 0.9637 0.9634
Accuracy 0.9642
Table 9: Classification metrics of the joint (ET+DECIGO) model in the SNR range 60–80.
Class Precision Recall $F_{1}$ score
PN 1.0000 1.0000 1.0000
UL 0.9997 1.0000 0.9999
SIS 1.0000 0.9914 0.9957
CIS 0.9915 1.0000 0.9957
NFW 1.0000 0.9998 0.9999
Macro average 0.9982 0.9982 0.9982
Accuracy 0.9983
Table 10: Classification metrics of the joint (ET+DECIGO) model in the SNR range 80–100.
Class Precision Recall $F_{1}$ score
PN 1.0000 1.0000 1.0000
UL 0.9997 1.0000 0.9999
SIS 0.9989 0.9522 0.9750
CIS 0.9544 1.0000 0.9767
NFW 1.0000 0.9988 0.9994
Macro average 0.9906 0.9902 0.9902
Accuracy 0.9903
At the lowest SNR interval (552020), model performance is strongly degraded by noise contamination. The accuracy reaches only $0.6339$, and the macro averaged $F_{1}$ score is $0.6059$ (Table 6). While the “PN” class is perfectly identified ($F_{1}=1.000$), the “UL” and three lensed classes exhibit substantial confusion. The confusion matrix (Fig. 12(a)) shows that SIS and CIS suffer the most from mutual misclassification, with correct recognition rates of about $39\%$ and $15\%$, respectively, and a fraction of samples incorrectly labelled as NFW. This indicates that at low SNR, waveform interference features are heavily masked by noise, hindering the network’s ability to extract stable phase dependent structures. The ROC curves (Figs. 11(d), 11(e), and 11(f)) show that the AUC values for the three lensed classes remain in the range 0.840.840.850.85, suggesting partial robustness to noise but limited discriminative confidence. In this regime, the joint model’s advantage is not yet fully realised, being primarily constrained by the low SNR boundary of the input data.
(a) micro average
(b) PN
(c) UL
(d) SIS
(e) CIS
(f) NFW
Figure 11: Comparison of ROC curves for five classification tasks across SNR intervals: (a) micro average, (b) PN, (c) UL, (d) SIS, (e) CIS, and (f) NFW.
As the SNR increases to moderate levels (20204040 and 40406060), the model performance improves markedly. In the 20204040 range, the accuracy rises to $0.7074$ (Table 7), and the macro averaged $F_{1}$ score reaches $0.6427$. Both SIS and NFW exhibit substantial gains relative to the low SNR regime, with $F_{1,\text{SIS}}$ improving from $0.4431$ to $0.5891$ and $F_{1,\text{NFW}}$ from $0.5496$ to $0.6620$. The CIS class, however, remains poorly classified, with $F_{1}=0.0215$.
At SNRs of 40406060, performance increases sharply: accuracy reaches $0.9642$ (Table 8), and the macro averaged $F_{1}$ rises to $0.9634$. The five classes become nearly perfectly separated, with “PN” and “UL” achieving $F_{1}=1.000$. Among the lensed signals, $F_{1,\text{SIS}}=0.9170$, $F_{1,\text{CIS}}=0.9009$, and $F_{1,\text{NFW}}=0.9995$. The confusion matrix (Fig. 12(c)) exhibits clear diagonal dominance, with residual misclassifications below $20\%$. The ROC curves (Fig. 11) show AUC values exceeding $0.9$ for all classes, with an average AUC near $0.99$, confirming the model’s capacity to distinguish diffraction and geometric magnification features at moderate SNRs. The accuracy, average $F_{1}$, and average AUC in Fig. 10 displays a distinct transition around an SNR of $40$, indicating that multiple detector data fusion becomes particularly effective beyond this threshold.
(a) SNR: 5–20
(b) SNR: 20–40
(c) SNR: 40–60
(d) SNR: 60–80
(e) SNR: 80–100
Figure 12: Confusion matrices of the joint (ET+DECIGO) model across different SNR intervals.
At high SNR (6060100100), model performance approaches saturation. In the 60608080 SNR interval, accuracy reaches $0.9983$ (Table 9), and the macro averaged $F_{1}$ is $0.9982$, with both Precision and Recall for all classes nearly unity. The confusion matrix (Fig. 12(d)) is almost perfectly diagonal, showing only negligible misclassifications.
In the 8080100100 SNR range, accuracy slightly decreases to $0.9903$ (Table 10), mainly due to minor boundary overlaps among classes, but remains close to ideal. The $F_{1}$ scores are $F_{\text{SIS}}=0.9750$, $F_{1,\text{CIS}}=0.9767$, and $F_{1,\text{NFW}}=0.9994$. The ROC curves (Figs. 11(a) and 11(b)) show $\mathrm{AUC}\approx 1.000$, confirming that the model achieves near perfect discrimination at high SNR. The confusion matrix (Fig. 12(e)) further verifies precise classification across all five classes, with virtually no off diagonal errors.
The overall trend (Fig. 10) reveals a rapid convergence of model performance with increasing SNR—from $63.4\%$ accuracy in the 552020 range to $99.8\%$ in the 60608080 interval—followed by stability at higher SNRs. These results confirm that the multiple detector configuration substantially enhances the extraction of interference and amplitude features at low to moderate SNRs, while maintaining calibrated probability outputs and precise classification at high SNR. Combining the results in Tables 610 and Figs. 1112, it is evident that the ET+DECIGO model achieves consistently high accuracy and robustness across all SNR regimes, highlighting the crucial role of multiband information fusion in detecting lensed GWs under noise conditions.
VII Conclusion and Outlook
This study proposes a systematic framework for lensed GWs identification under future multiband observations, integrating deep learning with multiband data fusion, and presents comprehensive experiments using simulated data from the ET and DECIGO detectors. The central goal is to develop and preliminarily evaluate an efficient algorithm capable of automatically identifying different lensing classes in complex noise environments, addressing the need for rapid and high precision classification in next generation detectors. The work first outlines the relevant physical mechanisms and then constructs a standardized multiple SNR dataset spanning an SNR range of $5$ to $100$ and five signal classes. The model employs an improved residual convolutional neural network that directly processes time domain waveforms. Two sets of experiments are conducted to evaluate model performance. The first trains single detector models for ET and DECIGO and a joint detector model for ET plus DECIGO across the full SNR range to evaluate the effect of combining detectors. The second trains the joint model within specific SNR intervals to analyze performance variation with signal strength. Results suggest that the joint model maintains stable accuracy at low SNR and achieves higher accuracy and macro averaged $F_{1}$ scores than either single detector model, indicating that multiband feature fusion may benefit weak signal recognition. The study provides a practical framework for multiband lensed GWs identification, demonstrates the capability of deep learning in complex noise environments, and offers a reproducible baseline for future joint observation data analysis.
Future research can further extend and refine this framework in several directions. In data generation, incorporating more diverse lens models, such as ellipsoidal potentials [hawken2009gravitational] and double lens systems [collett2016compound, vegetti2024strong], may enhance the adaptability and generalization of the model in complex lensing environments. On the algorithmic side, future studies could investigate network architectures that integrate time and frequency domain representations, or introduce attention mechanisms to better capture long timescale features [chen2023long, brauwers2021general]. Joint task learning strategies may also enable simultaneous lens classification and parameter estimation, for example, predicting lens mass or impact parameter. The proposed approach can be further evaluated under more realistic observational conditions, including multiple overlapping sources, non stationary noise, and anisotropic detector responses, to examine its robustness and scalability. As third generation ground based detectors (ET) and space detectors (DECIGO, LISA [amaro2017laser], Taiji [ruan2020taiji], TianQin [luo2016tianqin]) continue to advance, multiband and multiband observations are expected to play an increasingly important role in gravitational wave astronomy. Within this context, the present framework offers a useful foundation for rapid identification of lensed GWs, inference of lens parameters, and statistical investigations related to dark matter and cosmology, contributing to the ongoing development of precision gravitational wave astrophysics.
Acknowledgements
This work was supported by the National Key Research and Development Program of China (Grant No. 2021YFC2203004), the Fundamental Research Funds for the Central Universities Project (Grant No. 2024IAIS-ZD009), the National Natural Science Foundation of China (Grant Nos. 12575072 and 12347101), and the Natural Science Foundation of Chongqing (Grant No. CSTB2023NSCQ-MSX0103). This work made use of TensorFlow [Abadi_TensorFlow_Large-scale_machine_2015], PyCBC [alex_nitz_2024_10473621], GLoW_public [GLoW_public], and other open source Python packages.